OpenStudy (anonymous):

help check my answer! Point X is located at (2, 4), and Y is located at (−6, 10). Find the y value of the point that lies halfway between X and Y. 2 7 -3 -8 I said 7

OpenStudy (anonymous):

You could use midpoint formula \[\left( \frac{ x _{1}+x _{2} }{ 2 },\frac{ y _{1}+y _{2} }{ 2 }\right)\]\]

OpenStudy (anonymous):

thats what i used but im pretty sure i did it wrong

OpenStudy (anonymous):

No, you did it right :)

OpenStudy (anonymous):

so it is 7??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yeah 7 is right

OpenStudy (anonymous):

okay thank you guys so much!!